Find the equation of linear depreciation and the value of the car 5 years from now.
A car cost $20,000 and depreciates linearly at the rate of $3000 per year.

Answer by stanbon(75887)

Find the equation of linear depreciation and the value of the car 5 years from now.
A car cost $20,000 and depreciates linearly at the rate of $3000 per year.
-----
Value = 20,000-3000t
----
V(5) = 20,000-3000*5 = $5000.00
--------------
Cheers,
Stan H.
-------------